OpenStudy (anonymous):

Sam can paint a house in 5 days by himself, while Betty can paint the same house in 20 days. How long would it take Sam and Betty to paint the house together?

OpenStudy (anonymous):

4 days?? @superhelp101

OpenStudy (anonymous):

i doubt it

OpenStudy (anonymous):

should be \[\frac{5\times 20}{5+20}\] if that is 4 then you are right

OpenStudy (anonymous):

then reduce?

OpenStudy (anonymous):

yeah then compute

OpenStudy (anonymous):

its 4

OpenStudy (anonymous):

really ? ok then you win good work !

OpenStudy (anonymous):

100/25=4

OpenStudy (anonymous):

looks good to me not too hard was it?

OpenStudy (anonymous):

nope!